The heights of adult men in America are normally distributed, with a mean of 69.4 inches and a standard deviation of 2.64 inches. The heights of adult women in America are also normally distributed, but with a mean of 64.4 inches and a standard deviation of 2.57 inches.

a) If a man is 6 feet 3 inches tall, what is his z-score (to two decimal places)? z =

b) If a woman is 5 feet 11 inches tall, what is her z-score (to two decimal places)? z =

c) Who is relatively taller? The 5 foot 11 inch American woman The 6 foot 3 inch American man
Answer #1

Solution:-

For men

Mean = 69.4, S.D = 2.64

For Women

Mean = 64.4, S.D = 2.57

a) If a man is 6 feet 3 inches tall, what is his z-score is 2.12.

x = 6*12 + 3 = 75

By applying normal distribution:-

z = \frac{x-\mu }{\sigma }

z = 2.12

b) If a woman is 5 feet 11 inches tall, what is her z-score is 2.57.

x = 5*12 + 11 = 71

By applying normal distribution:-

z = \frac{x-\mu }{\sigma }

z = 2.57

c) The 5 foot 11 inch American woman is relatively taller because her z-score is greater than the -score for men.